OpenStudy (anonymous):

What is the multiplicative inverse of 2 1/2?

OpenStudy (solomonzelman):

what makes it into 1?

OpenStudy (solomonzelman):

\[5/2 \times x=1~~~~~~~~~~~find~~x.\]

OpenStudy (solomonzelman):

(HINT: eliminate x by dividing both sides by 5/2) Don't forget that \[\frac{x}{y} \div \frac{c}{d}=\frac{x}{y} \times \frac{d}{c}\]

OpenStudy (anonymous):

2/5?

OpenStudy (solomonzelman):

Yes!!!!!!!

OpenStudy (anonymous):

thanks!

OpenStudy (solomonzelman):

Anytime!